Last year, Silver Company's total variable production costs were $7,500, and its total fixed manufacturing overhead costs were $4,500. The company produced 3,000 units during the year and sold 2,400 units. There were no units in the beginning inventory. Which of the following statements is true?

  • Under variable costing, the average cost of the units in the ending inventory will be $4 each.

  • The operating income under absorption costing for the year will be $900 lower than the operating income under variable costing.

  • The ending inventory under variable costing will be $900 lower than the ending inventory under absorption costing.

  • Under absorption costing, the average cost of the units in ending inventory will be $2.50 each.
Answer #1

Answer :

C) The Ending inventory under Variable costing will be $900 Lower than the Ending inventory under Absorption costing

Variable costing

Unit Cost under Variable costing = 7,500/2,400 = 3.125

Ending Units = 3000-2400 =600

Ending inventory = 3.125*600 =1,875

Absorption Costing

Units cost Under Absorption Costing = 3.125+(4500/3000) =4.625

Ending invenotry = 4.625*600 = 2,775